# Adversarial Markov Games: On Adaptive Decision-Based Attacks and Defenses

### v2

@@ III Theoretical Framework
To investigate the robustness to evasion of real-world systems, two closely related perspectives are pertinent: a) thwarting decision-based attacks, and b) adapting attacks and evasive capabilities in tandem. For the latter, when attacks (and defenses) are evaluated in a non-adaptive manner, in the expansive sense we outlined in subsection II-C, results are rendered incomplete and untrustworthy [14, 25]. Nevertheless, when offensive or defensive techniques adapt the environments that they inhabit become non-stationary [30], putting further pressure on the IID foundations that ML builds on. To understand the implications of such techniques adapting, we perform a theoretical analysis of the possible interactions around the exposed interface of an ML-based system, interactions that can be more generally considered as sequential zero-sum games [31, 32, 27]. We initiate our analysis from adversaries, as the original factor and cause. In the following sections, introduced terms and notation are highlighted in orange.
@@ III-A Attacks
The most compelling threat that deployed ML-based systems face are decision-based black-box attacks, where no access is assumed to the model or its parameters, only the capacity to submit queries and receive hard-label responses. One of the first decision-based attacks was Boundary Attack [6]. A large number of others followed – each inventive in its own way – that manage to improve the overall performance, typically measured as the lowest perturbation achieved for the minimal amount of queries submitted. Prominent examples are HSJA [7], Guessing Smart (BAGS) [29], Sign-Opt [33], Policy-driven (PDA) [8], QEBA [34], and SurFree [35].
White-box attacks like C&W [17] cannot function in black-box environments as there is no closed-form description of the inference pipeline. To facilitate optimization, decision-based attacks commonly initialize from a sample belonging to the target class, as it can be considered an adversarial example with an unacceptably large perturbation. This switch allows the task to be solved continuously, by minimizing the perturbation while always staying on the adversarial side of the boundary. Decision-based attacks share further common aspects in their functioning, ones that we can abstract through: given starting and original samples $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}x_{g}$ and $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}x_{c}$ respectively, the goal is to iteratively propose adversarial candidates $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}x_{t}$, until the distance $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\delta=d(x_{t},x_{c})$ is minimized. This process follows different algorithmic approaches that represent different geometrical intuitions; we can describe it more generally by means of a candidate generation policy:
$\pi_{\theta}^{\mathcal{A}}=P\>(x_{t}|x_{g},x_{c},p^{\mathcal{A}},s^{\mathcal{A}})$ (2)
that given $x_{g}$ and $x_{c}$, with $p^{\mathcal{A}}$ the parameters and $s^{\mathcal{A}}$ the state of the attack, generates a candidate $x_{t}$. If we consider that attacks execute over discrete time steps, and assume the model always answers, then the attack procedure can be construed as a Markov Decision Process (MDP) to be solved, by finding the parameters $\theta$ that minimize $\delta$ in the least amount of queries.
Consider now a multinomial image classification model $\mathcal{M}$ under attack, with a discriminant function $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}F:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$, that for each input $x\in[0,1]^{d}$ generates an output $y:=\{y\in[0,1]^{m}|\sum_{c=1}^{m}y_{c}=1\}$ – a probability distribution over the $m$ classes. By definition, black-box environments preclude access to these probabilities; instead one can only observe the decision of the classifier $C$ that returns the highest probability class:
@@
$\psi(x_{t})=\begin{cases}1&\text{if}\quad C(x_{t})=c_{o}\\ -1&\text{if}\quad C(x_{t})\neq c_{o}\end{cases}$ (4)
As long as the model responds, $\psi$ can always be evaluated, it thus constitutes an essential mechanism upon which decision-based attacks build. The adversarial goal can then be described as the following constrained optimization problem:
@@
where the distance metric $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}d$ is an $\ell_{p}$-norm, with $p\in\{0,1,2,\inf\}$. As the threshold between adversarial and non-adversarial relies on the subjectivity of human perception, this optimization task highlights the fuzzy nature of adversarial examples, something that is further aggravated in domains where visual proximity is of little importance. Successful or unsuccessful adversarial examples are therefore delimited by an threshold $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\epsilon$ on perturbation, where $d(x,x_{t})\leq\epsilon$. While this constraint might not intuitively translate to non-visual domains, we nonetheless contend that minimal perturbation still remains a desideratum of adversarial examples.
The fact that real-world attacks are black-box and decision-based does not make them any less effective. For instance, HSJA is guaranteed to converge to a stationary point of Eq. (5). Given typical $\epsilon$ values for imperceptibility, this results in high attack success rates, even against adversarially trained models. The limited effectiveness of adversarial training against decision-based attacks can be attributed to the fundamentally out-of-distribution (OOD) nature of adversarial examples, and the inherent challenges in the saddle point optimization problem in Eq. (1) that make it difficult for algorithms to converge to a global solution. Additionally, decision-based attacks are challenging to incorporate during stochastic gradient descent (SGD) because these attacks rely on boundary navigation, and the farther the model is from convergence, the less effective the attack is. Scalability is another concern: while adversarial training typically involves a few steps (1-50) of white-box attacks like FGSM or PGD, decision-based attacks can take orders of magnitude more steps (queries) to produce an adversarial example.
@@
where $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}x^{i}_{b}$ is the best adversarial example generated by policy $\pi_{\theta}^{\mathcal{A}}$ during episode $i$. Given the dimensionality, it can be intractable to learn a policy that modifies the inputs at the feature space directly [36]; CIFAR-10, for instance, has more than 3K features to perturb.
In AI-enabled systems, the best practice is to freeze the model after validation so that no novel issues are introduced by retraining: for all queries $x_{t}$ submitted during an attack session, we can therefore assume that $F_{0}=F_{1}=...=F_{t},\forall t$. While this is representative of real-world settings, it is also what enables adversaries to discover adversarial examples that were not identified beforehand. The fact however that adversaries follow a candidate generation policy, introduces a querying behavior which can be observed and exploited by a defensive methodology. Consequently, while model-hardening approaches like adversarial training are necessary, they can be insufficient in defending from decision-based attacks.
@@
Let $F_{c}$ be the discriminant function of the adversarially trained model $\mathcal{M}$. Then in order for $\mathbb{E}[\sum_{i=1}^{N}d(x^{i}_{b},x^{i}_{c})]\geq\epsilon$ in HSJA, two capabilities are necessary: a) a decision function $D^{\prime}\neq\operatorname*{arg\,max}$, and b) additional context $\tau$ s.t for adversarial query $x_{t}$, $C(x_{t})=D(F_{c}(x_{t}))\neq D^{\prime}(\tau,F_{c}(x_{t}))$.
Intuitively, HSJA operates in 3 stages which repeat: a binary search that puts $x_{t}$ on the decision boundary, a gradient estimation step, and projection step along the estimated gradient. If the model always responds truthfully, the adversary will be able to accurately perform all these steps and converge to the optimal adversarial; without loss of generality, we can extend this intuition to other decision-based attacks which navigate the boundary. Secondly, the model should be able to differentiate between two, otherwise identical, queries when one is part of an attack and the other is not, and this is possible through a stateful representation; see Appendix A for the proof.
@@ III-B Defenses
Proposition III.1 indicates that alternative classification policies are necessary in the presence of decision-based attacks, for instance classification with rejection or intentional misdirection. In related work, rejection has been realized in the form of conformal prediction where model predictions are sets of classes including the empty one, or learning with rejection [11, 37]; while misdirection has emerged as a technique in adversarial RL and cybersecurity domains [10, 12]. While adversarially training the discriminant function $F$ empirically shows the capacity to resist decision-based attacks, the manner in which the model responds has a complementary potential. This gap between the empirical and theoretically possible robustness to decision-based attacks is the locus where a distinct from model hardening, active and adaptive defense can emerge. Active defenses have direct implications on attacks themselves however. Let us now assume an agent carrying out an active defense policy:
@@
with $x_{t}$ the query, $s^{\mathcal{D}}_{t}$ the state for the defense as created by past queries, and $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\alpha$ the binary decision: when the query is deemed adversarial, it is rejected by returning $\alpha=1$. If this policy is stationary, the environment dynamics in turn become stationary too, and thus together with the adversarial task itself, bypassing the defense can also be formulated as an MDP to be solved. In two-player, zero-sum games, the moment an agent follows a stationary policy, it becomes exploitable through the reward obtained by an adversary [38]. Active defenses, a consequence of decision-based attacks, entail therefore adaptive adversaries.
@@
A proof for BAGS and HSJA is included in Appendix A. An adversary can reason, as a corollary to Proposition III.1, that such defenses have to be in place as it is suboptimal not too. However, there is a second reason to consider adaptive attacks even in the absence of active defenses, as attack policies are often suboptimal with their default, empirically defined parameters. Adapting attack policies is essentially the optimization of these parameters, and as an approach has proven very effective in other black-box or expensive-to-evaluate contexts, like Neural Architecture Search and Data Augmentation [39, 40, 41]. Our results in Section V further indicate the correspondence between adaptive and self-optimizing, showing that adaptive attacks consistently outperform non-adaptive, particularly against active defenses.
Consider now an active defense that is based on a similarity or a conformal metric. In the twofold meaning we introduced in Section II, adaptive attack implies the capability to bypass a similarity based defense; adaptive control implies optimization instead, the active tuning of all the available tools to evade the defense and minimize the perturbation (cf. Fig. 1). The revised adversarial objective then is to find the optimal policy that also evades detection, and the way to achieve this is by adapting the candidate generation policy (2) itself. Notably, and despite the black-box and discontinuous nature of the task, this optimization can be fully gradient-based. Decision-based attacks can recover gradient-based solutions to their objective, despite neither the active defense nor the model itself being accessible in closed-form. For model $\mathcal{M}$, adversarial queries $x_{t}$, and active defense $\pi_{\phi}^{\mathcal{D}}$ making decisions $\alpha_{t}$, we can thus formulate the following:
@@
The proof is included in Appendix A. So far we have established that, a) in the presence of decision-based attacks, active defenses are necessary, yet conditional on adversarial agency they are insufficient and, b) adaptive attacks can become optimal in terms of both evasion and efficiency by observing and adapting to the discrete model decisions. To complete the puzzle, the last piece is turning active defenses also adaptive.
@@
Figure 2: Schematic model of an AMG environment. Due to the inherent uncertainty of behavior at either side of the interface, it is a partially observable RL environment mirrored for each agent where one’s decisions become the other’s observations. (I) denotes an adaptive attacker (cf. Fig. 1), (II) model hardening (passive defense), and (III) an active defense.
@@
As offensive and defensive policies are strictly competitive, we can define a reward $\rho$ of the defensive policy as $\rho(\tau_{i})=\sum_{x_{t}\in\tau_{i}}\alpha_{t}$. Then by fixing the optimal evasive policy $\pi_{\theta^{*}}^{\mathcal{A}}$, and updating $\pi_{\phi}^{\mathcal{D}}$ with expected reward $\mathbb{E}_{\pi^{\mathcal{D}}_{\phi}}[\rho(\tau_{i})]$ gradients, at convergence the resulting ${\pi^{\mathcal{A}}_{\theta^{*}}}$ will be the optimal defensive policy. ∎
@@ III-C Adversarial Markov Games
By reasoning on both offensive and defensive capabilities, we highlighted why we cannot consider them independently. As adaptive attacks and defenses are logical consequences of each other, their composition results in a turn-taking competitive game. A precise game-theoretic formulation requires however the exact analytical description of the environment: the model, the players and their utility functions, as well as the permitted interactions and the transition dynamics, something exceedingly intractable in this as well as most other cybersecurity contexts. Model-free methods however can learn optimal offensive and defensive responses directly through interaction with their environment [12, 42], obviating the need to learn a model of it or to find exact solutions to the bi-level optimization problem in Eq. (1) that is NP-hard [43].
To that end, Turn-Taking Partially-Observable Markov Games (TT-POMGs) introduced by Greenwald et al. [44] is a generalization of Extensive-Form Games (EFGs), widely used representations for non-cooperative, sequential decision-making games of imperfect and/or incomplete information. TT-POMGs are an apt formalism for decision-based attacks and defenses, with the additional useful property that they can be transformed to equivalent belief state MDPs, significantly simplifying their solution.
The competition underlying adversarial example generation has been explored in no-box and white-box settings [27, 45]. We instead focus on decision-based, interactive environments, and assume that they have unknown but stationary dynamics: any agency present is considered part of the environment and therefore fixed in its behavior. By folding the strategies of other agents into the transition probabilities and the initial probability distribution of the game, an optimal policy computed in the resulting MDP will correspond to the best-response strategy in the original TT-POMG. The congruence between TT-POMGs and MDPs is useful not only from a theoretical perspective, but also for its practical implications in the security of ML-based systems: provided that adversarial agents and their capabilities can be identified through rigorous threat modeling, computing the best response strategy in the simulated environment will correspond to the optimal defense.
The environment that encompasses adversarial attacks, adversarial defenses, and benign queries, we describe it as an Adversarial Markov Game (AMG) – a special case of TT-POMG – and is depicted in Figure 2. Formally, we represent AMG as a tuple $\langle i,S,O,A,\tau,r,\gamma\rangle$
@@
We have shown that an adaptive defense policy $\pi^{\mathcal{D}}_{\phi}$ is necessary to deter decision-based attacks, and that consequently the candidate generation policy $\pi^{\mathcal{A}}_{\theta}$ has to be also be adaptive. As with plausible assumptions we cannot assume access to the exact state of the other agent, the states $O^{\mathcal{D}}$, $O^{\mathcal{A}}$ are partial observations of the complete state $\mathcal{S}$ of the full game. When the competing agents that hold beliefs about each other are human, they engage in recursive reasoning expressed as [I believe that [my opponent believes [that I believe…]]]. In the study of opponent modeling, considering other agent policies as stationary and part of the environment is equivalent to 0th level recursive reasoning: the agent models how the opponent behaves based on the observed history, but not how the opponent would behave based on how the agent behaves [46, 47]. In this work we consider more involved recursive reasoning out of scope, as AMGs can be solved by single-agent RL algorithms, and perform the empirical evaluation without building explicit models of opponent behavior.
IV Environment Specification

### v3

@@ III Theoretical Framework
In this section, we abstract through the individual properties of decision-based attacks and defenses to extract more general insights than a purely empirical study would render. To investigate how robust real-world systems are to evasion, two related perspectives are crucial: a) resisting decision-based attacks, and b) adapting attacks and evasive capabilities together. When attacks (and defenses) are evaluated in a non-adaptive manner, in the expansive sense we outlined in subsection II-C, results are unreliable [14, 26]. Note that with offensive or defensive methodologies adapting, their environments become non-stationary [31], putting further pressure on the IID foundations that ML builds on. To understand the implications of this adaptation, we perform an analysis of the possible interactions on the interface of an ML-based system, interactions that can be more generally considered as sequential zero-sum games [32, 33, 29]. In the following sections, introduced terms and notation are highlighted in orange.
@@ III-A Attacks
The most compelling threat that deployed ML-based systems face are decision-based black-box attacks, where no access is assumed to the model or its parameters, only the capacity to submit queries and receive hard-label responses. One of the first decision-based attacks was Boundary Attack [6], followed by others that improve the overall performance, typically measured as the lowest perturbation achieved for the minimal amount of queries submitted. Prominent examples are HSJA [7], Guessing Smart (BAGS) [28], Sign-Opt [34], Policy-driven (PDA) [8], QEBA [35], and SurFree [36].
White-box attacks like C&W [20] do not function in black-box environments, as there is no access to the inference pipeline. To facilitate optimization, decision-based attacks commonly initialize from a sample belonging to the target class, as it can be considered an adversarial example with an unacceptably large perturbation. This switch allows the task to be solved continuously, by minimizing the perturbation while always staying on the adversarial side of the boundary. Decision-based attacks share further common aspects in their function that we can abstract through: given starting and original samples $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}x_{g}$ and $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}x_{c}$ respectively, the goal is to iteratively propose adversarial candidates $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}x_{t}$, until the distance $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\delta=d(x_{t},x_{c})$ is minimized. This process follows different algorithmic approaches that represent different geometrical intuitions; we can describe it more generally by means of a candidate generation policy:
$\pi_{\theta}^{\mathcal{A}}=P\>(x_{t}|x_{g},x_{c},p^{\mathcal{A}},s^{\mathcal{A}}),$ (2)
that given $x_{g}$ and $x_{c}$, with $p^{\mathcal{A}}$ the parameters and $s^{\mathcal{A}}$ the state of the attack, generates a candidate $x_{t}$. As attacks execute over discrete time steps, if we assume that the model always answers the attack procedure can be construed as a Markov Decision Process (MDP) to be solved, by finding the parameters $\theta$ that minimize $\delta$ for a given number of queries.
Consider now a multinomial image classification model $\mathcal{M}$ under attack, with a discriminant function $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}F:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$, that for each input $x\in[0,1]^{d}$ generates an output $y:=\{y\in[0,1]^{m}|\sum_{c=1}^{m}y_{c}=1\}$ – a probability distribution over the $m$ classes. As black-box environments prevent access to these probabilities, one can only observe the decision of the classifier $C$ that returns the highest probability class:
@@
$\psi(x_{t})=\begin{cases}+1&\text{if}\quad C(x_{t})=c_{o}\\ -1&\text{if}\quad C(x_{t})\neq c_{o}\end{cases}$ (4)
As long as the model responds, $\psi$ can always be evaluated, and constitutes the fundamental mechanism upon which decision-based attacks build. The adversarial goal can then be described as the following constrained optimization problem:
@@
where the distance metric $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}d$ is an $\ell_{p}$-norm, with $p\in\{0,1,2,\inf\}$. As the threshold between adversarial and non-adversarial relies on the subjectivity of human perception, this highlights the indefinite nature of adversarial examples, further exemplified in domains where visual proximity is of little importance. Successful or unsuccessful adversarial examples are therefore delimited by an threshold $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\epsilon$ on perturbation, where $d(x,x_{t})\leq\epsilon$.
Real-world attacks being black-box does not make them less effective. For instance, HSJA is guaranteed to converge to a stationary point of Eq. (5). Given typical $\epsilon$ values for imperceptibility, this results in high attack success rates, even against adversarially trained models. The limitations of adversarial training against decision-based attacks can be attributed to the out-of-distribution (OOD) nature of adversarial examples, and the saddle point optimization problem of Eq. (1) that make it difficult for algorithms to converge to a global solution. Furthermore, incorporate decision-based attacks in training is not scalable as it can take orders of magnitude more steps (queries) to produce an adversarial example, than white-box attacks which take a few steps (1-50 in e.g. PGD [4]).
@@
where $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}x^{i}_{b}$ is the best adversarial example generated by policy $\pi_{\theta}^{\mathcal{A}}$ during episode $i$. Given the dimensionality of the input, it can be intractable to learn a policy that modifies the feature space directly [37]; CIFAR-10, for instance, has more than 3K features to perturb.
In AI-based systems, the best practice is to freeze the model after validation so that no novel issues are introduced by retraining: for all queries $x_{t}$ submitted during an attack session, we can therefore assume that $F_{0}=F_{1}=...=F_{t},\forall t$. While this is representative of real-world settings, it also enables adversaries to discover adversarial examples that were not identified beforehand. Consequently, while model-hardening through adversarial training is necessary, it can also be insufficient against decision-based attacks like HSJA.
@@
Let $F_{c}$ denote the discriminant function of an adversarially trained model $\mathcal{M}$, and let $C(x)=D(F_{c}(x))$ denote its classifier. Then in HSJA, to satisfy $\mathbb{E}[\sum_{i=1}^{N}d(x_{b}^{i},x_{c}^{i})]\geq\epsilon$ it is necessary that: (a) $D\neq\operatorname*{arg\,max}$, and (b) context $\tau$ exists s.t. for some query $x_{t}$, $D(F_{c}(x_{t}))\neq D^{\prime}(\tau,F_{c}(x_{t}))$, where $D^{\prime}$ is a stateful extension of $D$.
Intuitively, HSJA operates in 3 stages which repeat: a binary search that puts $x_{t}$ on the decision boundary, a gradient estimation step, and projection step along the estimated gradient. If the model always responds truthfully, the adversary will be able to accurately perform all these steps and converge to the optimal adversarial; without loss of generality, we can extend this intuition to other decision-based attacks which navigate the boundary. Secondly, the model should be able to distinguish between two, otherwise identical, queries, when one is part of an attack and the other is not, a capability achievable through statefulness; see Appendix A for the proof.
@@ III-B Defenses
Proposition III.1 suggests that alternative classification policies are necessary in the presence of decision-based attacks, e.g. classification with rejection or intentional misdirection. Rejection has been realized in the form of conformal prediction, where model predictions are sets of classes including the empty one, or learning with rejection [11, 38]; while misdirection has emerged as a technique in adversarial RL and cybersecurity domains [10, 12]. While adversarially training the discriminant function $F$ empirically shows some degree of robustness to decision-based attacks, the manner in which the model responds has a complementary potential. The gap between the empirical and theoretically achievable robustness is the source for an active defense distinct from model hardening. Active defenses have direct implications on attacks themselves however. Let us now assume an agent carrying out an active defense policy:
@@
with $x_{t}$ the query, $s^{\mathcal{D}}_{t}$ the state for the defense as created by past queries, and $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\boldsymbol{\alpha}$ the binary decision: for queries deemed adversarial, $\boldsymbol{\alpha}=1$, otherwise $\boldsymbol{\alpha}=0$. When this policy is stationary, the environment dynamics become stationary in turn, thus besides the adversarial task itself, bypassing the defense can also be formulated as an MDP to be solved (Figure 2). In two-player, zero-sum games, the moment an agent follows a stationary policy, it becomes exploitable through the reward obtained by an adversary [39]. Active defenses, as consequence of decision-based attacks, entail therefore adaptive adversaries.
@@
A proof for BAGS and HSJA is included in Appendix A. An adversary can reason, as a corollary to Proposition III.1, that such defenses have to be in place as it is suboptimal not too. However, there is a second reason to consider adaptive attacks even in the absence of active defenses, as attack policies are often suboptimal with their default, empirically defined parameters. Adapting attack policies is essentially the optimization of these parameters, and as an approach has proven very effective in other black-box or expensive-to-evaluate domains, like Neural Architecture Search and Data Augmentation [40, 41, 42]. Our results in section V further indicate the correspondence between adaptive and self-optimizing, showing that adaptive consistently outperform non-adaptive attacks, particularly against active defenses.
Consider now an active defense that is based on a similarity or conformal metric. In the twofold meaning we introduced in subsection II-C, adaptive attack implies the capability to bypass a similarity based defense; adaptive control implies optimization instead, the active tuning of all the available tools to evade the defense and minimize the perturbation (cf. Figure 1). The updated adversarial objective then is to find the optimal policy that also evades detection, and the way to achieve this is by adapting the candidate generation policy (2) itself. Notably, and despite the black-box and discontinuous nature of the task, this optimization can be fully gradient-based. Decision-based attacks can recover gradient-based solutions to their objective, despite neither the active defense nor the model itself being accessible in closed-form. For model $\mathcal{M}$, adversarial queries $x_{t}$, and active defense $\pi_{\phi}^{\mathcal{D}}$ making decisions $\alpha_{t}$, we can thus formulate the following:
@@
The proof is included in Appendix A. We thus have established that, a) in the presence of decision-based attacks, active defenses are necessary, yet conditional on adversarial agency they are insufficient and, b) adaptive attacks can become optimal in terms of both evasion and efficiency by observing and adapting to the discrete model decisions. To complete the puzzle, the last piece is turning active defenses also adaptive.
@@
Since the game is zero-sum, we may define the defensive policy reward $\rho$ on any trajectory $\tau_{i}=(x_{1},\dots,x_{T})$ as $\rho(\tau_{i})=\sum_{x_{t}\in\tau_{i}}\alpha_{t}$. Treating the adversary’s policy $\pi_{\theta^{*}}^{\mathcal{A}}$ as fixed, we perform gradient ascent on the expected reward $J(\phi)\;=\;\mathbb{E}_{\tau_{i}\sim(\pi_{\theta^{*}}^{\mathcal{A}},\,\pi_{\phi}^{\mathcal{D}})}\bigl{[}\rho(\tau_{i})\bigr{]}$. Under standard smoothness assumptions, this converges to $\phi^{*}=\arg\max_{\phi}J(\phi)$, which is precisely the defender’s best response to $\pi_{\theta^{*}}^{\mathcal{A}}$. ∎
@@ III-C Adversarial Markov Games
By reasoning on both offensive and defensive capabilities, we highlight why one cannot consider them independently. As adaptive attacks and defenses are logical consequences of each other, their composition forms a turn-taking competitive game. A precise game-theoretic formulation requires full knowledge of the environment: its models, players and their utility functions, as well as the permitted interactions and the transition dynamics, something typically intractable in this and other cybersecurity settings. Model-free methods however can learn optimal (offensive and defensive) responses directly through interaction with their environment [12, 43], avoiding the need for explicit modeling or solving the NP-hard bi-level optimization problem of Eq. (1[44].
To that end, Turn-Taking Partially-Observable Markov Games (TT-POMGs), introduced by Greenwald et al. [45], generalize Extensive-Form Games (EFGs) that model non-cooperative, sequential decision-making games of imperfect and/or incomplete information. TT-POMG is a suitable formalism for decision-based attacks and defenses, with the added benefit that it can be transformed into an equivalent belief state MDP, significantly simplifying its solution.
Prior work has explored the competition underlying adversarial example generation in no-box and white-box settings [29, 46]. We instead focus on decision-based, interactive environments, with unknown but stationary dynamics: all other agents present are considered part of the environment and therefore fixed in their behavior. By folding the strategies of other agents into the transition probabilities and the initial probability distribution of the game, an optimal policy computed in the resulting MDP will correspond to the best-response strategy in the original TT-POMG. The congruence between TT-POMGs and MDPs has both theoretical and practical value for securing AI-based systems: once adversarial agents and their capabilities are identified through rigorous threat modeling, the best-response strategy in the simulated environment yields the optimal defense.
Figure 2: Schematic model of an AMG environment. Due to the inherent uncertainty of behavior at either side of the interface, it is a partially observable MDP, mirrored for each agent where one’s decisions become the other’s observations. (I) denotes an adaptive attacker (cf. Fig. 1), (II) model hardening (passive defense), and (III) an active defense.
We describe the environment that encompasses adversarial attacks, adversarial defenses, and benign queries, as an Adversarial Markov Game (AMG) – a special case of TT-POMG – depicted in Figure 2. Formally, we represent AMG as a tuple $\langle i,S,O,A,\tau,r,\gamma\rangle$
@@
We have shown that an adaptive defense policy $\pi^{\mathcal{D}}_{\phi}$ is necessary to deter decision-based attacks, and that consequently the candidate generation policy $\pi^{\mathcal{A}}_{\theta}$ has to be also be adaptive. As with plausible assumptions we cannot assume access to the exact state of the other agent, the states $O^{\mathcal{D}}$, $O^{\mathcal{A}}$ are partial observations of the complete state $\mathcal{S}$ of the full game. For instance, when the competing agents (holding beliefs about each other) are human, they engage in recursive reasoning expressed as [I believe that [my opponent believes [that I believe…]]]. In the study of opponent modeling, considering other agent policies as a stationary part of the environment is equivalent to 0th level recursive reasoning: the agent models how the opponent behaves based on the observed history, but not how the opponent would act based on how the agent behaves [47, 48]. In this work we consider more involved recursive reasoning out of scope, as AMGs can be solved by single-agent RL algorithms, and perform the empirical evaluation without building explicit models of opponent behavior.
IV Threat Model